### v1

# Born-Infeld Electrogravity and Dyonic Black Holes

Abstract
Born-Infeld electrogravity is defined through a Lagrangian that couples gravity and electromagnetism within a single determinantal structure. The field equations are derived in Palatini’s formalism, where the metric, connection, and vector potential are varied independently in the action. As a result, the gravitational sector reduces to Einstein’s equations with a torsion-free, metric-compatible connection. The electrodynamic sector, in turn, admits two equivalent interpretations or pictures: it can be seen either as a standard Born-Infeld electrodynamics in an effective background geometry, or as an anomalous Born-Infeld electrodynamics in the physical metric. We illustrate the dynamics by analyzing the horizon structure and extremality conditions of spherically symmetric dyonic solutions. A comparison with the Reissner-Nordström geometry shows that Born-Infeld electrogravity softens but does not eliminate curvature divergences, and geodesic incompleteness persists.
I Introduction
Born-Infeld theory (BI) has its origins in the context of nonlinear electrodynamics. Proposed by Max Born and Leopold Infeld in 1934 [1, 2, 3, 4], this theory emerged as an extension of classical Maxwell electromagnetism, with the aim of removing the divergences in the self-energy of the point-like charged particle by the following action
$S\propto\int d^{4}x\sqrt{-\det\left(g_{\mu\nu}+bF_{\mu\nu}\right)}-\lambda\sqrt{-\det(g_{\mu\nu})},$ (1)
where $g_{\mu\nu}$ is the metric tensor, $F_{\mu\nu}$ is the electromagnetic field tensor, $b$ is a new fundamental constant which sets the order of magnitude where the nonlinear effects become relevant, and $\lambda$ controls the asymptotic behavior of vacuum solutions, with $\lambda=1$ corresponding to Minkowski spacetime. In the weak field limit, the BI theory reduces to the Maxwell action plus small corrections, while for the strong regime the field equations deviate significantly from Maxwell’s electromagnetism. This well known capacity to cure the singularities motivated the so-called BI gravity theories, where the gravitational action is built by analogy with the BI structure (Eq. (1)) through the Ricci tensor $R_{\mu\nu}$
$S\propto\int d^{4}x\sqrt{-\det\left(g_{\mu\nu}+bR_{\mu\nu}\right)}-\lambda\sqrt{-\det(g_{\mu\nu})}.$ (2)
This action was studied in the standard metric formalism [5, 6, 7] as well as in the Palatini (metric-affine) approach [8] with the intention of regularizing the spacetime singularities inside black holes and also in the initial singularity at the beginning of the Universe in the Big Bang model [9]. For a recent review on BI inspired modifications of gravity, see [10].
Many modified gravity theories inspired by the BI action introduce matter fields following the standard prescription of minimal coupling in metric theories of gravity (MTG) [11]. This approach ensures consistency with Einstein’s equivalence principle [12], and splits the total action into a gravitational part and a matter part: $S_{\text{total}}=S_{\text{gravity}}+S_{\text{matter}}$. The matter sector is then constructed from the Minkowskian metric promoted to a curved spacetime metric $g_{\mu\nu}$, making use of covariant derivatives $\nabla_{\mu}$, and some minimal coupling prescription (which is not always free of ambiguities [13]). Explicitly, one writes $S_{\text{matter}}=S[g_{\mu\nu},\psi_{m},\nabla\psi_{m}]$, where $\psi_{m}$ generically denotes the matter fields. Perhaps for this reason, attempts to unify the matter and gravity sectors within a single determinantal action have received relatively little attention. To the best of our knowledge, the notable exception is the class of models proposed by Vollick in [14], who introduced an action of the form
$S\propto\int d^{4}x\sqrt{-\det\left(g_{\mu\nu}+\epsilon R_{\mu\nu}(\Gamma)+\beta M_{\mu\nu}\right)}-\lambda\sqrt{-\det(g_{\mu\nu})},$ (3)
where $M_{\mu\nu}$ denotes a tensor constructed from matter fields, and $\epsilon$ and $\beta$ (the inverse of the Born-Infeld constant $b$) are coupling constants. For massive scalar fields one may take $M_{\mu\nu}=\partial_{\mu}\phi\partial_{\nu}\phi$, while for electromagnetic fields $M_{\mu\nu}=F_{\mu\nu}$. A construction of this type does not fit into the usual separation of matter and gravity in theories based on the principle of minimal coupling, which suggests that such models should eventually conflict with the experimental evidence supporting the equivalence principle. Nevertheless, Vollick showed that for $M_{\mu\nu}=\partial_{\mu}\phi\partial_{\nu}\phi$, the field equations coincide with those of GR minimally coupled to a free scalar field. For electromagnetic and fermionic fields, at the first non-trivial order in perturbations, one recovers the usual Einstein-Maxwell and Einstein-Dirac systems, indicating that such theories effectively reduce to an MTG representation at that order.
The main purpose of this work is to investigate whether Born-Infeld electrogravity corrections can smooth or avoid the inherent singularities of GR solutions, while keeping the asymptotic behavior unchanged. For that purpose, we employ the gravito-electromagnetic Lagrangian by following the pioneering Vollick proposal (Eq. (3)), recently revisited by Afonso et al [15], with $M_{\mu\nu}=F_{\mu\nu}$ and we deduce the field equation by varying the action à la Palatini (see, for instance, [16]), where the connection and the metric are treated as independent variables. In particular, we analyze here a dyonic black hole, a spherically symmetric spacetime with both electric and magnetic charges, and find that the singularity is partially smoothed.
The dynamical equations for the gravitational field reduce to Einstein’s equations with a torsion-free, metric-compatible connection. In turn, the electromagnetic field equations admit two equivalent interpretations or pictures: (i) a standard Born-Infeld electrodynamics in an effective background geometry, or (ii) an anomalous Born-Infeld electrodynamics in the physical metric. This BI electrogravity ends up being equivalent to General Relativity sourced by the anomalous Born-Infeld field. Suggestively, this resembles the correspondence between modified gravity and General Relativity plus different matter fields [17, 18]. To the best of our knowledge, the formalism for relating both pictures is presented here for the first time.
The article is organized as follows. Section II introduces the gravito-electromagnetic Lagrangian with non-minimal coupling (Eq. (3)), and we obtain the field equations in the Palatini formalism. Section III exploits the structure of Born–Infeld-like Lagrangians to obtain the dynamics. Section IV shows that, despite appearances, the system evolves consistently with an MTG theory. Section V exhibits the results concerning the dyonic black hole, which carries both electric and magnetic charges. The horizons and singularities are analyzed; in the asymptotic weak field regime, the solution is shown to approach the Reissner–Nordström geometry. Finally, the conclusions are presented in Section VI.
II Born-Infeld electrogravity
II.1 Born-Infeld electrodynamics
Born-Infeld Lagrangian was initially introduced as an idea to avoid the divergence of the electrostatic field of a point-like charge [1, 2, 3]. In its most primitive version, it simply mimicked the Lagrangian of relativistic mechanics, since its electrostatic form was
$L\propto\sqrt{1-b^{-2}\mathbf{E}^{2}}\ .$ (4)
This was a successful attempt, in the sense that the electric field $\mathbf{E}$ of a point-like charge turned out to be bounded above by the new Born-Infeld constant $b$, in the same way that $c$ is an upper limit for speed in relativistic mechanics. Born and Infeld soon realized that this Lagrangian was the Minkowskian version of a more general one, whose geometric formulation in terms of the electromagnetic field $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ is given by the density [4]
$L_{BI}[A]=-\frac{b^{2}}{4\pi}\;\left(\sqrt{|\det(g_{\mu\nu}+b^{-1}F_{\mu\nu})|}-\sqrt{|\det(g_{\mu\nu})|}\right)\ .$ (5)
[…]
Born-Infeld electrodynamics can be summarized through its dynamical equation and energy-momentum tensor, [19, 20, 21, 22]
$\partial_{\mu}\left(\sqrt{-g}\;\mathcal{F}^{\mu\nu}\right)=0\ ,$ (6)
$T^{\mu\nu}_{BI}\equiv-\frac{2}{\sqrt{-g}}\;\frac{\delta L_{BI}}{\delta g_{\mu\nu}}=\frac{1}{4\pi}\left[{\mathcal{F}}^{\mu\rho}\,F^{\;\nu}_{\rho}-b^{2}\;g^{\mu\nu}\left(1-\sqrt{1+b^{-2}2S-b^{-4}P^{2}}\right)\right]\ ,$ (7)
(metric signature ++---; Gaussian units) where $g\equiv\det(g_{\mu\nu})$, and $\mathcal{F}^{\mu\nu}$ is
$\mathcal{F}^{\mu\nu}\equiv\frac{F^{\mu\nu}-b^{-2}P\ {}^{\ast}F{{}^{\mu\nu}}}{\sqrt{1+b^{-2}2S-b^{-4}P^{2}}}\ .$ (8)
II.2 Born-Infeld electrogravity
Inspired by Born-Infeld nonlinear electrodynamics, and motivated by its capabilities of smoothing singularities, we will study an electrogravity theory governed by the Lagrangian [14, 15]
$L\propto\sqrt{-q}-\lambda\sqrt{-g}\ ,$ (9)
where $q\equiv\det(q_{\mu\nu})$ is the determinant of an auxiliary tensor, defined as
$q_{\mu\nu}\equiv g_{\mu\nu}+\epsilon R_{(\mu\nu)}(\Gamma)+\beta F_{\mu\nu}(A).$ (10)
$F_{\mu\nu}(A)=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ is the electromagnetic field tensor, and $R_{(\mu\nu)}(\Gamma)$ is the symmetric part of the Ricci tensor, constructed from a metric-independent affine connection $\Gamma^{\lambda}_{\mu\nu}$, given by  [^2] We use the convention $V^{\mu}_{\ ;\nu}=V^{\mu}_{\ ,\nu}+\Gamma^{\mu}_{\rho\nu}V^{\rho}$.
$R_{\mu\nu}(\Gamma)\equiv\partial_{\alpha}\Gamma^{\alpha}_{\mu\nu}-\partial_{\nu}\Gamma^{\alpha}_{\mu\alpha}+\Gamma^{\beta}_{\mu\nu}\Gamma^{\alpha}_{\beta\alpha}-\Gamma^{\beta}_{\mu\alpha}\Gamma^{\alpha}_{\beta\nu}.$ (11)
The parameters $\epsilon$ and $\beta$ are fundamental constants of the theory, with dimensions of squared length and inverse field strength, respectively. $\epsilon^{-1}\beta^{2}$ is proportional to the Newton constant; $\lambda\neq 1$ indicates the presence of a cosmological constant $\Lambda=\epsilon^{-1}(1-\lambda)$, as will be shown.
II.3 Dynamical equations à la Palatini
In general, the variation of the determinant of a matrix $M$ is
$\delta\;\ln|\det M|=\operatorname{Tr}({\bar{M}}\cdot\delta M)\ ,$ (12)
where $\bar{M}$ denotes the inverse of M{M}. Thus, the variation of the Lagrangian (9) yields
$\delta L\propto\sqrt{-q}\,{\bar{q}}^{\nu\mu}\,\delta q_{\mu\nu}-\lambda\sqrt{-g}\,g^{\mu\nu}\,\delta g_{\mu\nu}\ ,$ (13)
where $\bar{q}^{\mu\nu}$ is the inverse of $q_{\mu\nu}$,
$q_{\mu\alpha}\,{\bar{q}}^{\alpha\nu}=\delta_{\mu}^{\nu}\ ,$ (14)
and
$\delta q_{\mu\nu}=\delta g_{\mu\nu}+\epsilon\,\delta R_{(\mu\nu)}(\Gamma)+\beta\,\delta F_{\mu\nu}({A})\ .$ (15)
In the Palatini formalism [16], $\bf g$}, $\bf\Gamma$}, and $\bf A$} are independent dynamical variables. The variation of the Lagrangian with respect to each of them leads to three equations governing the coupled dynamics of the metric, the affine connection, and the electromagnetic field.
The equation resulting from the variation with respect to the metric,
$\sqrt{-q}\,\bar{q}^{(\mu\nu)}-\lambda\sqrt{-g}\,g^{\mu\nu}=0\,$, (16)
is not strictly a dynamical equation since it does not contain second-order time derivatives; actually it is a constraint equation, as it relates the canonical variables and their momenta. This is a consequence of the absence of metric derivatives in the Lagrangian. Instead the Lagrangian contains derivatives of $\bf\Gamma$} and $\bf A$}; so the dynamics will follow from varying it with respect to these variables.
The variation with respect to $\bf A$} leads to a dynamical equation for the electromagnetic field in a given geometric background. Since $\delta F_{\mu\nu}=2\,\partial_{[\mu}\delta A_{\nu]}$, after integrating by parts one obtains
$\partial_{\mu}\!\left(\sqrt{-q}\,\bar{q}^{[\mu\nu]}\right)=0\,.$ (17)
So, the symmetric and antisymmetric parts of the tensor density $\sqrt{-q}\;\bar{q}^{\mu\nu}$ separately enter the constraint equation (16) and the dynamical equation (17), as a nice consequence of the symmetric and antisymmetric characters of $\delta g_{\mu\nu}$ and $\delta F_{\mu\nu}$ respectively.
On the other hand, the variation of the Ricci tensor in Eq. (11) with respect to the affine connection $\bf\Gamma$} can be written as
$\delta R_{\mu\nu}=\nabla^{\Gamma}_{\alpha}(\delta\Gamma^{\alpha}_{\mu\nu})-\nabla^{\Gamma}_{\nu}(\delta\Gamma^{\alpha}_{\mu\alpha})+T^{\theta}_{\ \lambda\nu}\,\delta\Gamma^{\lambda}_{\mu\theta}\ ,$ (18)
where $T^{\theta}_{\ \alpha\beta}\equiv 2\Gamma^{\theta}_{[\beta\alpha]}$ is the torsion tensor. In this expression for $\delta R_{\mu\nu}$ we take advantage of the fact that the difference $\delta{\bf\Gamma}$ between two affine connections is a tensor (even though $\bf\Gamma$} is not).
Notice that the symmetric character attributed to $R_{\mu\nu}$ forces the variation $\delta{\bf\Gamma}$ to ensure that $\delta R_{[\mu\nu]}=0$; therefore, the variation $\delta{\bf\Gamma}$ cannot be considered arbitrary. However, the projective-type contribution [23, 24] to $\delta R_{[\mu\nu]}$ coming from $\delta\Gamma^{\alpha}_{\mu\alpha}$ in the second term of Eq. (18) should not worry; its contribution to the variation of the Lagrangian is degenerate with the one coming from $\delta A_{\mu}$ and can be absorbed (see Appendix B). Actually, condition $\delta R_{[\mu\nu]}=0$ can be enforced by just fixing the torsion zero; in this way it is $\delta\Gamma^{\alpha}_{[\mu\nu]}=0$ (because the torsion is fixed), and no contributions to $\delta R_{[\mu\nu]}$ from the first and third terms in Eq. (18) will apprear. So Eq. (18) becomes
[…]
and the corresponding variation of the Lagrangian yields the dynamical equation  [^3] As follows from the integration by parts leading to this equation, $\nabla^{\Gamma}_{\alpha}\left(\sqrt{-q}\,\xi^{\nu\mu}\right)$ denotes the tensor density $\sqrt{-q}\,\nabla^{\Gamma}_{\alpha}\xi^{\nu\mu}+\xi^{\nu\mu}\,\partial_{\alpha}\sqrt{-q}$.
$\nabla^{\Gamma}_{\alpha}\left(\sqrt{-q}\,{\bar{q}}^{(\nu\mu)}\right)=0\ .$ (20)
However, the constraint (16) turns this equation into the condition for the connection being metric-compatible. Combined with the torsion-free condition imposed above, this uniquely determines the connection to be the Levi-Civita connection. Thus, Eqs. (16) become dynamic, since it now contains second derivatives of the metric. Equations (16) and (17) govern the dynamics of the metric and electromagnetic fields. To make practical use of this system of equations, one must either explicitly construct the tensor $\bar{q}^{\mu\nu}$, or find a shortcut to avoid that task. In this sense, we may notice that from the electrodynamical point of view, the Lagrangian (9)–(10) corresponds to a Born-Infeld electrodynamics in a geometry where the symmetric tensor
$\mathcal{G}_{\mu\nu}\equiv g_{\mu\nu}+\epsilon R_{(\mu\nu)}$ (21)
plays the role of the background metric. This means that Eq. (17) must correspond to a Born-Infeld equation in the “metric” $\mathcal{G}_{\mu\nu}$. On the other hand, the variation of the Lagrangian (9)–(10) with respect to the metric $g_{\mu\nu}$ should be related to the energy-momentum tensor of the Born-Infeld field in the background geometry $\mathcal{G}_{\mu\nu}$. Since Born-Infeld electrodynamics in arbitrary geometric backgrounds is well known, these elements could help to gain insight into several aspects of Born-Infeld electrogravity, before undertaking the task of building the tensor $\bar{q}^{\mu\nu}$.
III Dynamics of Born-Infeld electrogravity
III.1 The inverse of $q_{\mu\nu}$
Any matrix $q$ can be decomposed into the sum of its symmetric and antisymmetric parts, $\mathcal{G}$ and $F$,
$q=\mathcal{G}+F=\mathcal{G}(I+\bar{\mathcal{G}}\ F)\equiv\mathcal{G}(I+\breve{F})\ ,$ (22)
[…]
$(I+\breve{F})^{-1}=\sum_{n=0}^{\infty}(-\breve{F})^{n}~.$ (23)
But in four dimensions there exists another antisymmetric matrix which helps us to express $(I+\breve{F})^{-1}$ in terms of $F$ without resorting to a series: the dual matrix of $F$ whose components are  [^5] The “metric” role of $\mathcal{G}$ also implies that $^{\ast}\breve{F}_{\mu\nu}=\frac{\sqrt{-\mathcal{G}}}{2}~\epsilon_{\lambda\rho\mu\nu}~{\breve{F}}^{\lambda\rho}=\frac{1}{2}~\breve{\epsilon}_{\lambda\rho\mu\nu}~{\breve{F}}^{\lambda\rho}$, where ${\breve{F}}^{\lambda\rho}=\bar{\mathcal{G}}^{\lambda\mu}F_{\mu\nu}\bar{\mathcal{G}}^{\nu\rho}$. Besides, ${{}^{\ast}}{{}^{\ast}}\equiv-1$.
$^{\ast}\breve{F}^{\mu\nu}\equiv\frac{1}{2\,\sqrt{-\mathcal{G}}}~\epsilon^{\lambda\rho\mu\nu}~F_{\lambda\rho}=\frac{1}{2}\ \breve{\varepsilon}^{\lambda\rho\mu\nu}~F_{\lambda\rho}~,$ (24)
where $\epsilon^{\lambda\rho\mu\nu}$ is the Levi-Civita symbol ($\epsilon^{0123}=1=\epsilon_{0123}$), and $\breve{\varepsilon}^{\lambda\rho\mu\nu}$ is the Levi-Civita tensor associated with the “metric” $\mathcal{G}$.
We will operate with the matrices $\breve{F}$ and $^{\ast}\breve{F}$ of components  [^6] The product of matrices is the contraction of inner indices: $\bar{q}\,q=I$ means $\bar{q}^{\mu\lambda}q_{\lambda\nu}=\delta^{\mu}_{\nu}$.
$\breve{F}_{~\nu}^{\mu}=\bar{\mathcal{G}}^{\mu\lambda}F_{\lambda\nu}\ ,\ \ \ \ \ \ ^{\ast}\breve{F}_{~\nu}^{\mu}={{}^{\ast}\breve{F}}^{\mu\lambda}~\mathcal{G}_{\lambda\nu}~.$ (25)
$\breve{F}$ and $^{\ast}\breve{F}$ are the blocks to build two matrices proportional to the identity:
$2\breve{S}~I$ $=$ $-\breve{F}\breve{F}+(^{\ast}{\breve{F}})~^{\ast}\breve{F}~\,$, (26)
$\breve{P}~I$ $=$ $-\breve{F}\,~{}^{\ast}\breve{F}=-(^{\ast}{\breve{F}})\ \breve{F}\ ,$ (27)
where the scalar $\breve{S}$ and the pseudoescalar $\breve{P}$ are obtained by tracing the former equations:
$\breve{S}$ $=-\frac{1}{4}~\breve{F}_{~\lambda}^{\mu}\ \breve{F}_{~\mu}^{\lambda}=\frac{1}{4}~{}^{\ast}\breve{F}_{~\lambda}^{\mu}\,{}^{\ast}\breve{F}_{~\mu}^{\lambda}~,$ (28)
$\breve{P}$ $=-\frac{1}{4}{}~\breve{F}_{~\lambda}^{\mu}\,{}^{\ast}\breve{F}_{~\mu}^{\lambda}=-\frac{1}{4}{}~{F}_{\mu\lambda}\,^{\ast}\breve{F}^{\lambda\mu}~.$ (29)
Now, let us compute
$(I+\breve{F})~(I-\breve{F}+{{}^{\ast}\breve{F}}\,^{\ast}\breve{F}+a~^{\ast}\breve{F})$~ $=$ $I-\breve{F}~\breve{F}+{{}^{\ast}\breve{F}}~^{\ast}\breve{F}+a~^{\ast}\breve{F}+\breve{F}~^{\ast}\breve{F}~^{\ast}\breve{F}+a~\breve{F}~^{\ast}\breve{F}$ (30)
$=$ $I+2\breve{S}~I+a~^{\ast}\breve{F}-\breve{P}~I~^{\ast}\breve{F}-a~\breve{P}~I$
By choosing $a$ equal to $\breve{P}$, one has the result that
$(I+\breve{F})^{-1}=\frac{I-\breve{F}+{{}^{\ast}\breve{F}}~^{\ast}\breve{F}+\breve{P}~^{\ast}\breve{F}}{1+2\breve{S}-\breve{P}^{2}}$ (31)
So the inverse of $q=\mathcal{G}+\beta~F=\mathcal{G}(I+\beta~\breve{F})$ is
$\bar{q}=\frac{I-\beta~\breve{F}+\beta^{2}~{}^{\ast}\breve{F}~^{\ast}\breve{F}+\beta^{3}~\breve{P}~^{\ast}\breve{F}}{1+2\beta^{2}~\breve{S}-\beta^{4}~\breve{P}^{2}}~\mathcal{\bar{G}}\ .$ (32)
By combining this expression for $\bar{q}$ with the value for $\det{q}$ computed in Appendix A (Eq. (104)), one obtains
$\sqrt{-q}\ \bar{q}^{\mu\nu}=\sqrt{-\mathcal{G}}~\frac{\mathcal{\bar{G}}^{\mu\nu}-\beta~\breve{F}^{\mu\nu}+\beta^{2}~{}^{\ast}\breve{F}_{~\lambda}^{\mu}~{}^{\ast}\breve{F}^{\lambda\nu}+\beta^{3}~\breve{P}~^{\ast}\breve{F}^{\mu\nu}}{\sqrt{1+2\beta^{2}~\breve{S}-\beta^{4}~\breve{P}^{2}}}~.$ (33)
We will end this preparatory subsection by noting that
$\sqrt{-\mathcal{G}}\ ^{\ast}\breve{F}^{\mu\lambda}=\frac{1}{2}\ \epsilon^{\eta\rho\mu\lambda}~F_{\eta\rho}=\sqrt{-g}\ ^{\ast}{F}^{\mu\lambda}~;$ (34)
according to the relation (27) this means that
$\sqrt{-\mathcal{G}}\ \breve{P}=\sqrt{-g}\ P\ .$ (35)
On the other hand, by multiplying Eq. (26) with $^{\ast}\breve{F}$, and using Eq. (27), one gets
$2\breve{S}~^{\ast}\breve{F}=\breve{P}~\breve{F}+(^{\ast}{\breve{F}})~(^{\ast}{\breve{F}})~^{\ast}\breve{F}\ .$ (36)
III.2 Electrodynamics
In Eq. (17) let us replace $\sqrt{-q}\;\bar{q}^{[\mu\nu]}$ with the antisymmetric part of Eq. (33); then
$\partial_{\mu}\left(\sqrt{-\mathcal{G}}\ \frac{\breve{F}^{\mu\nu}-\beta^{2}\ \breve{P}\ ^{\ast}\breve{F}^{\mu\nu}}{\sqrt{1+2\beta^{2}\ \breve{S}-\beta^{4}\ \breve{P}^{2}}}\right)=0\ .$ (37)
As anticipated, this equation governs the standard Born-Infeld electrodynamics in the geometric background described by the “metric” $\mathcal{G}_{\mu\nu}$, since we recognize in parentheses the respective tensor $\breve{\mathcal{F}}^{\mu\nu}$ as introduced in Eq. (8). Certainly, we do not yet know $\mathcal{G}_{\mu\nu}=g_{\mu\nu}+\epsilon R_{\mu\nu}$ and its inverse $\bar{\mathcal{G}}^{\mu\nu}$, since they are subject to the gravitodynamic equations.
III.3 Gravitodynamics
Now, replace $\sqrt{-q}\;\bar{q}^{(\mu\nu)}$ in Eq. (16) with the symmetric part of Eq. (33); then
$\sqrt{-\mathcal{G}}~\frac{\mathcal{\bar{G}}^{\mu\nu}+\beta^{2}~{}^{\ast}\breve{F}_{~\lambda}^{\mu}~{}^{\ast}\breve{F}^{\lambda\nu}}{\sqrt{1+2\beta^{2}~\breve{S}-\beta^{4}~\breve{P}^{2}}}=\lambda\sqrt{-g}\,g^{\mu\nu}\ .$ (38)
After contracting this expression with $\mathcal{G}_{\nu\rho}$, and substituting
$\sqrt{-\mathcal{G}}=\sqrt{-g}\;\sqrt{\det(\delta^{\mu}_{\nu}+\epsilon R^{\mu}_{\;\nu})}\ ,$ (39)
where $R^{\mu}_{\;\nu}=g^{\mu\rho}R_{\rho\nu}$, the dynamical equations for the geometry become
$\lambda\,\frac{\delta^{\mu}_{\nu}+\epsilon\,R^{\mu}_{\ \nu}}{\sqrt{\det(\delta^{\mu}_{\nu}+\epsilon\,R^{\mu}_{\ \nu})}}=\frac{\delta^{\mu}_{\nu}+\beta^{2}~{}^{\ast}\breve{F}_{~\lambda}^{\mu}~{}^{\ast}\breve{F}^{\lambda}_{\ \nu}}{\sqrt{1+2\beta^{2}~\breve{S}-\beta^{4}~\breve{P}^{2}}}\ .$ (40)
Ricci tensor both in Eqs. (37) and (40) is written in terms of the Levi-Civita connection, as obtained from the variational calculus in Section (II.3).
We remark that any geometry whose Ricci tensor is $R^{\mu}_{\;\nu}=-\Lambda\delta^{\mu}_{\nu}$ –in particular, de Sitter geometry– is a vacuum solution to Eq. (40) for $\lambda=1-\epsilon\Lambda$. [^7] For $^{\ast}\breve{F}^{\mu}_{\ \lambda}=0$ the equation reduces to $\lambda\,\frac{(1-\epsilon\Lambda)}{\sqrt{(1-\epsilon\lambda)^{4}}}\,\delta^{\mu}_{\nu}=\delta^{\mu}_{\nu}$.
IV Equivalence principle
Equations (37) and (40) seem to imply that the equivalence principle is violated by the action (9). This impression is caused by the presence of the Ricci tensor $R_{\mu\nu}$ in Eq. (37) which governs the dynamics of the electromagnetic field. In fact, the Ricci tensor plays a role in the volume $\sqrt{-\mathcal{G}}$ and the tensor $\breve{F}^{\mu\nu}=\bar{\mathcal{G}}^{\mu\lambda}\bar{\mathcal{G}}^{\nu\rho}F_{\lambda\rho}$. In addition, the source in the r.h.s of Eq. (40) is also contaminated by the Ricci tensor. However, as shown in Ref. [15], the dynamical equations can be combined in such a way that both undesired contributions of the Ricci tensor disappear. We will use the rest of this subsection to demonstrate this property.
Let us begin with Eq. (38), which is the result of varying the Lagrangian (9) with respect to the metric $g_{\mu\nu}$. This dynamical equation can be regarded as the relation between the operations of raising indices with $\bar{\mathcal{G}}^{\mu\nu}$ or $g^{\mu\nu}$. Although the term $^{\ast}F\ ^{\ast}F$ on the l.h.s of Eq. (38) is annoying, its presence can be easily handled with the help of Eqs. (27) and (36), depending on whether $^{\ast}F\ ^{\ast}F$ acts on $F$ or $^{\ast}F$. For instance, by contracting Eq. (38) with $F_{\nu\rho}$ and using Eq. (27) one obtains
$\sqrt{-\mathcal{G}}~\frac{\breve{F}^{\mu}_{\ \rho}-\beta^{2}\;\breve{P}\ ^{\ast}\breve{F}_{~\rho}^{\mu}}{\sqrt{1+2\beta^{2}~\breve{S}-\beta^{4}~\breve{P}^{2}}}=\lambda\sqrt{-g}\ F^{\mu}_{\ \rho}\ ,$ (41)
which expresses the relationship that dynamics establishes between electrodynamic variables generated by the “metric” $\mathcal{G}$ and those associated with the metric $g$. To further explore this relationship, we raise the index $\rho$ in Eq. (41) by contracting again with Eq. (38); thus it results that
$\mathcal{G}~\frac{(\breve{F}^{\mu}_{\ \rho}-\beta^{2}\;\breve{P}\ ^{\ast}\breve{F}_{~\rho}^{\mu})~(\mathcal{\bar{G}}^{\rho\nu}+\beta^{2}~{}^{\ast}\breve{F}_{~\lambda}^{\rho}~{}^{\ast}\breve{F}^{\lambda\nu})}{1+2\beta^{2}~\breve{S}-\beta^{4}~\breve{P}^{2}}=\lambda^{2}g\ F^{\mu\rho}\ ,$ (42)
which, via Eq. (36) is converted into
$\mathcal{G}~\frac{(1+\beta^{4}\breve{P}^{2})~\breve{F}^{\mu\nu}-2\beta^{2}\;\breve{P}~(1+\beta^{2}\breve{S})\ ^{\ast}\breve{F}^{\mu\nu}}{1+2\beta^{2}~\breve{S}-\beta^{4}~\breve{P}^{2}}=\lambda^{2}g\ F^{\mu\nu}\ .$ (43)
We can also find the relationship between dual matrices. For this, we contract the former equation with the Levi-Civita symbol $\epsilon_{\mu\nu\lambda\rho}$, which must be associated with factors $\sqrt{-\mathcal{G}}$ and $\sqrt{-g}$ already present in the l.h.s. and the r.h.s., respectively:
$\sqrt{\mathcal{G}}~~\frac{(1+\beta^{4}\breve{P}^{2})~^{\ast}\breve{F}_{\lambda\rho}+2\beta^{2}\;\breve{P}~(1+\beta^{2}\breve{S})\ \breve{F}_{\lambda\rho}}{1+2\beta^{2}~\breve{S}-\beta^{4}~\breve{P}^{2}}=\lambda^{2}\sqrt{g}~~^{\ast}F_{\lambda\rho}\ .$ (44)
We are ready to get the relationship between the field invariants. By contracting Eq. (43) with Eq. (44), and then taking the trace, the result is
$(-\mathcal{G})^{3/2}~\frac{(1+\beta^{4}\breve{P}^{2})^{2}-4\beta^{4}\;\breve{P}^{2}~(1+\beta^{2}\breve{S})^{2}+4\beta^{2}\ \breve{S}\ (1+\beta^{4}\breve{P}^{2})(1+\beta^{2}\breve{S})}{(1+2\beta^{2}~\breve{S}-\beta^{4}~\breve{P}^{2})^{2}}\ \breve{P}=\lambda^{4}(-g)^{3/2}\ P\ ,$ (45)
which simplifies to yield
$(-\mathcal{G})^{3/2}~\breve{P}=\lambda^{4}(-g)^{3/2}\ P\ ,$ (46)
and together with Eq. (35) leads to the dynamical relationships between the pseudoscalars and the volumes,
$\lambda^{2}\ \breve{P}=P\ ,\ \ \ \ \ \ \ \ \sqrt{-\mathcal{G}}=\lambda^{2}\sqrt{-g}\ .$ (47)
In this way, the general relation (34) is dynamically converted into
$\lambda^{2}\ {}^{\ast}\breve{F}^{\mu\lambda}=\ ^{\ast}{F}^{\mu\lambda}~.$ (48)
On the other hand, by contracting Eq. (43) with $F_{\mu\nu}$, then taking the trace and using $\mathcal{G}=\lambda^{4}\ g$, one obtains the expression of $S$ in terms of $\breve{S}$ and $\breve{P}$:
$\lambda^{2}~\frac{\breve{S}\ (1-\beta^{4}\;\breve{P}^{2})-2\ \beta^{2}\ \breve{P}^{2}}{1+2\beta^{2}~\breve{S}-\beta^{4}~\breve{P}^{2}}=S\ .$ (49)
Now, let us focus on a calculation concerning the structure of Born-Infeld field equation (8). Eqs. (48), (43) and (47) can be combined to yield
${F}^{\mu\nu}+\beta^{2}\lambda^{-2}P\;^{\ast}{F}^{\mu\nu}$ $=$ $\lambda^{2}\frac{(1+\beta^{4}\breve{P}^{2})\breve{F}^{\mu\nu}-2\beta^{2}\breve{P}(1+\beta^{2}\breve{S})^{\ast}\breve{F}^{\mu\nu}}{1+2\beta^{2}~\breve{S}-\beta^{4}~\breve{P}^{2}}+\lambda^{2}\ \beta^{2}\breve{P}\ ^{\ast}\breve{F}^{\mu\nu}$ (50)
$=$ $\lambda^{2}\frac{(1+\beta^{4}\breve{P}^{2})~(\breve{F}^{\mu\nu}-\beta^{2}\breve{P}\ ^{\ast}\breve{F}^{\mu\nu})}{1+2\beta^{2}~\breve{S}-\beta^{4}~\breve{P}^{2}}\ .$
Notice that Eq. (49) implies
$1-2(\beta/\lambda)^{2}\;S-(\beta/\lambda)^{4}\;P^{2}=1-2\beta^{2}\ \left(\frac{\breve{S}\ (1-\beta^{4}\;\breve{P}^{2})-2\ \beta^{2}\ \breve{P}^{2}}{1+2\beta^{2}~\breve{S}-\beta^{4}~\breve{P}^{2}}\right)-\beta^{4}\breve{P}^{2}=\frac{(1+b^{-4}\breve{P}^{2})^{2}}{1+2\beta^{2}~\breve{S}-\beta^{4}~\breve{P}^{2}}\ ,$ (51)
which allows Eq. (50) to be written as
$\lambda^{-2}\ \frac{{F}^{\mu\nu}+(\beta/\lambda)^{2}\;P\ ^{\ast}{F}^{\mu\nu}}{\sqrt{1-2(\beta/\lambda)^{2}\;S-(\beta/\lambda)^{4}\;P^{2}}}=\frac{\breve{F}^{\mu\nu}-\beta^{2}\breve{P}\ ^{\ast}\breve{F}^{\mu\nu}}{\sqrt{1+2\beta^{2}~\breve{S}-\beta^{4}~\breve{P}^{2}}}\ .$ (52)
This equation can be multiplied by volume $\sqrt{-\mathcal{G}}$ (so $\sqrt{-g}$ will appear in the l.h.s., according to Eq. (47)) to get on the right side the density tensor that enters the electrodynamic equation (37).
Equation (52) is a direct consequence of applying the dynamical equation (38). It has a profound meaning; the electrodynamics that emerges from Lagrangian (9) can be understood both as the standard Born-Infeld electrodynamics in the geometric background of “metric” $\mathcal{G}_{\mu\nu}=g_{\mu\nu}+\epsilon R_{\mu\nu}$ (where the electromagnetic field seems to be coupled to the curvature), or as an anomalous Born-Infeld electrodynamics in the geometric background of metric $g_{\mu\nu}$,
$\partial_{\mu}\left(\sqrt{-{g}}\ \frac{{F}^{\mu\nu}+(\beta/\lambda)^{2}\;P\ ^{\ast}{F}^{\mu\nu}}{\sqrt{1-2(\beta/\lambda)^{2}\;S-(\beta/\lambda)^{4}\;P^{2}}}\right)=0\ .$ (53)
We distinguish these two possibilities by calling them picture $\mathcal{G}$ and picture $g$ respectively. We say that the picture $g$ is anomalous because $S$ appears with the “wrong” sign in the l.h.s. of Eq. (52) . So, for passing from the (standard) r.h.s. to the (anomalous) l.h.s. in Eq. (52) not only the metric must be changed, but $\beta^{2}$ must be replaced with $(\beta/\lambda)^{2}$ (or, alternatively, $\beta^{2}\rightarrow-\beta^{2}$, $F_{\mu\nu}\rightarrow\lambda^{-1}F_{\mu\nu}$). However, the picture $g$ is not intrinsically anomalous, since we could have started from a Lagrangian built with an imaginary value for $\beta$; in this way the anomalous character would be transferred to the picture $\mathcal{G}$, while the l.h.s of Eq. (52) would look as the standard Born-Infeld electrodynamics for the field $\lambda^{-1}F_{\mu\nu}$. Imaginary values for $\beta$ are not forbidden in this theory, since just even powers of $\beta$ appear in the dynamical equations (the Lagrangian (9) is made of even powers of $\beta$ according to Eq. (104) ).
Let us now focus on the gravitodynamic equation (40), where the Ricci tensor of the geometry $g_{\mu\nu}$ is sourced by an energy-momentum built in the picture $\mathcal{G}$, which implies that the curvature is on both sides of the equation. So we wonder whether it is possible to give this equation a form in picture $g$, where the curvature is sourced by an energy momentum that does not depend on the curvature. First, note that $\det(\delta^{\mu}_{\nu}+\epsilon\,R^{\mu}_{\ \nu})=\mathcal{G}/g$ is dynamically equal to $\lambda^{4}$ (see Eq. (47)), so it will be easy to rewrite Eq. (40) in terms of the Einstein tensor. In particular, the curvature scalar can be computed by taking the trace to the equation:
$\lambda^{-1}\,(4+\epsilon\,R)=\frac{4+\beta^{2}~4\breve{S}}{\sqrt{1+2\beta^{2}~\breve{S}-\beta^{4}~\breve{P}^{2}}}\ .$ (54)
Therefore, by using (26) and other properties already used, one gets
$\lambda^{-1}\epsilon\;(G^{\mu}_{\nu}-\Lambda\;\delta^{\mu}_{\nu})=\lambda^{-1}\epsilon\;(R^{\mu}_{\nu}-\frac{1}{2}\;R\;\delta^{\mu}_{\nu})-(\lambda^{-1}-1)\;\delta^{\mu}_{\nu}=\delta^{\mu}_{\nu}-\frac{\delta^{\mu}_{\nu}-\beta^{2}\ \breve{F}^{\mu\lambda}\;F_{\lambda\nu}}{\sqrt{1+2\beta^{2}~\breve{S}-\beta^{4}~\breve{P}^{2}}}\ .$ (55)
The r.h.s. is a necessarily conserved tensor. But it is not the electromagnetic energy-momentum tensor in picture $\mathcal{G}$ (nor is the l.h.s. of the equation the Einstein tensor of that picture). Therefore, we will try to give the r.h.s. the form of the energy-momentum electromagnetic tensor in the picture $g$. The energy-momentum (7) of standard Born-Infeld electrodynamics can be rewritten, among other equivalent ways, as
$4\pi\beta^{2}\ T^{\mu}_{\nu}=-\delta^{\mu}_{\nu}+\frac{\delta^{\mu}_{\nu}+\beta^{2}~{}^{\ast}{F}^{\mu\lambda}\;{}^{\ast}{F}_{\lambda\nu}}{\sqrt{1+2\beta^{2}~{S}-\beta^{4}~{P}^{2}}}\ .$ (56)
For the anomalous case, it then results in
$4\pi(\beta/\lambda)^{2}\ {T^{anom}}^{\mu}_{\;\nu}=-\delta^{\mu}_{\nu}+\frac{\delta^{\mu}_{\nu}-(\beta/\lambda)^{2}~{}^{\ast}{F}^{\mu\lambda}\;{}^{\ast}{F}_{\lambda\nu}}{\sqrt{1-2(\beta/\lambda)^{2}~{S}-(\beta/\lambda)^{4}~{P}^{2}}}\ ,$ (57)
which can be translated into the variables of picture $\mathcal{G}$ by means of (44) , (48) and (27),
$\delta_{\mu}^{\nu}-(\beta/\lambda)^{2}\ {}^{\ast}{F}^{\mu\lambda}\;{}^{\ast}{F}_{\lambda\nu}$ $=\delta_{\mu}^{\nu}-\beta^{2}\ \frac{(1+\beta^{4}\breve{P}^{2})~^{\ast}\breve{F}^{\mu\lambda}\;{}^{\ast}\breve{F}_{\lambda\nu}-2\beta^{2}\;\breve{P}^{2}~(1+\beta^{2}\breve{S})\ \delta^{\mu}_{\nu}}{1+2\beta^{2}~\breve{S}-\beta^{4}~\breve{P}^{2}}$
$=\delta^{\mu}_{\nu}\ \frac{1+2\beta^{2}\breve{S}-\beta^{4}\breve{P}^{2}+2\beta^{4}\breve{P}^{2}(1+\beta^{2}\breve{S})}{1+2\beta^{2}~\breve{S}-\beta^{4}~\breve{P}^{2}}-\frac{(1+\beta^{4}\breve{P}^{2})~\beta^{2}~{}^{\ast}\breve{F}^{\mu\lambda}\;{}^{\ast}\breve{F}_{\lambda\nu}}{1+2\beta^{2}~\breve{S}-\beta^{4}~\breve{P}^{2}}$
$=\frac{1+\beta^{4}\breve{P}^{2}}{1+2\beta^{2}~\breve{S}-\beta^{4}~\breve{P}^{2}}\ \left((1+2\beta^{2}\breve{S})\ \delta^{\mu}_{\nu}-\beta^{2}~{}^{\ast}\breve{F}^{\mu\lambda}\;{}^{\ast}\breve{F}_{\lambda\nu}\right)$
$=\frac{1+\beta^{4}\breve{P}^{2}}{1+2\beta^{2}~\breve{S}-\beta^{4}~\breve{P}^{2}}\ \left(\delta^{\mu}_{\nu}-\beta^{2}~\breve{F}^{\mu\lambda}\;{F}_{\lambda\nu}\right)\ .$ (58)
So, by applying Eq. (51) it results that
$\frac{\delta_{\mu}^{\nu}-(\beta/\lambda)^{2}\ {}^{\ast}{F}^{\mu\lambda}\;{}^{\ast}{F}_{\lambda\nu}}{\sqrt{1-2(\beta/\lambda)^{2}~{S}-(\beta/\lambda)^{4}~{P}^{2}}}=\frac{\delta^{\mu}_{\nu}-\beta^{2}~\breve{F}^{\mu\lambda}\;{F}_{\lambda\nu}}{\sqrt{1+2\beta^{2}~\breve{S}-\beta^{4}~\breve{P}^{2}}}\ .$ (59)
Thus, the source in Eq. (55) is the anomalous energy-momentum tensor of picture $g$:
$G^{\mu}_{\nu}-\Lambda\;\delta^{\mu}_{\nu}=\frac{4\pi\beta^{2}}{\epsilon\lambda}\ {T^{anom}}^{\mu}_{\;\nu}\ .$ (60)
As already stated, Born-Infeld electrogravity depends only on the even powers of $\beta$. In case an imaginary value for $\beta$ is chosen –to avoid the anomalous behavior in the picture gg– the sign of $\epsilon$ must be consistently changed to preserve the positive sign of Newton’s constant in the above equation (cf. [25]). In such a case, the cosmological constant in the former equation must be understood as $(1-\lambda)/(-\epsilon)$.
V Dyonic Black Holes
Equations (53) and (60) make up an Einstein-Born-Infeld system that is equivalent to the system formed by Eqs. (37) and (40). Its spherically symmetric solutions have been widely studied, and we refer the reader to [10] (and references therein) for a comprehensive review on this topic. This geometry has been studied in standard Einstein-Born-Infeld contexts [26, 27] and also in the anomalous determinantal context [15].
We are interested in the capabilities of this type of theory for smoothing singularities; as an example we solve the field equations for the solution corresponding to a dyonic black hole, where both the electric and magnetic monopoles are present.
V.1 Electrodynamics
To work with Eq. (53), we propose for the electromagnetic field the closed 2-form
$\textbf{F}=-e(r)\ dt\wedge dr-g(\theta)\ d\theta\wedge d\phi\ ,$ (61)
in a spherically symmetric geometric background given by
$g_{\mu\nu}=\text{diag}\left(f(r),-f^{-1}(r),-r{{}^{2}},-r{{}^{2}}\sin{{}^{2}}\theta\right)\ .$ (62)
$f(r)$ is a function to be determined by the geometrodynamics, but it decouples from the electrodynamical problem. In fact, $f(r)$ does not enter into the volume $\sqrt{-g}$ nor into the operation of raising indices to $F_{\mu\nu}$ in Eq.  (61). The field configuration (61) describes an electrostatic field $E_{r}=e(r)$ and a magnetostatic field $B_{r}=r^{-2}g(\theta)\sin\theta$, both in the radial direction. The spherical symmetry imposes that $g(\theta)=p\sin\theta$; thus the magnetic field is Coulombian,
$B_{r}=\frac{p}{r^{2}}\ .$ (63)
By replacing the field invariants $S=(p^{2}r^{-4}-e(r)^{2})/2$ and $P=-r^{-2}p\ e(r)$, Eq. (53) is integrated to yield
$e(r)=\frac{q}{\sqrt{r^{4}-(\beta/\lambda)^{2}(q^{2}+p^{2})}}\,.$ (64)
If $\beta$ is real, then $e(r)$ is real for $r>r_{o}\equiv\left[\beta^{2}\lambda^{-2}(q^{2}+p^{2})\right]^{1/4}$, due to the anomalous behavior in the picture $g$. The magnetic field (63) is regular in the domain $r\in(r_{o},\infty)$, but $e(r)$ diverges at $r=r_{o}$, Instead, if $\beta$ is imaginary the solution is defined in the domain $r\in(0,\infty)$; the electric field is regular everywhere (it is the standard Born-Infeld field of a point-like charge), but the magnetic field diverges at $r=0$.
The energy-momentum tensor (57) is
${T^{anom}}^{\mu}_{\;\nu}=\frac{\lambda^{2}}{4\pi\ \beta^{2}}\ \text{diag}\left(1-\sqrt{1-\frac{r_{o}^{4}}{r^{4}}}\ ,\ 1-\sqrt{1-\frac{r_{o}^{4}}{r^{4}}}\ ,\ 1-\frac{1}{\sqrt{1-\frac{r_{o}^{4}}{r^{4}}}}\ ,\ 1-\frac{1}{\sqrt{1-\frac{r_{o}^{4}}{r^{4}}}}\right)\ ,$ (65)
whose trace determines the behavior of the curvature scalar $R$ in Eq. (60):
[…]
Note that even though the electric and magnetic monopoles $q$ and $p$ enter the field invariants $S$ and $P$ in an unequal way, they contribute to the energy and pressure on an equal footing through the combination $\sqrt{q^{2}+p^{2}}$. This feature will be inherited by the geometry we are going to examine.
V.2 Gravitodynamics
The Ricci tensor $R^{\mu}_{\ \nu}$ for the metric (62) is
$R^{\mu}_{\ \nu}=\text{diag}\left(\frac{2f^{\prime}+rf^{\prime\prime}}{2r},\frac{2f^{\prime}+rf^{\prime\prime}}{2r},-\frac{1-f(r)-rf^{\prime}}{r^{2}},-\frac{1-f(r)-rf^{\prime}}{r^{2}}\right),$ (67)
Then, to solve the gravitodynamic equation (60) with the source (65), $f$ must be a solution to the first order equation
$\left(\frac{r}{r_{o}}\right)^{2}+\lambda\ \sqrt{\left(\frac{r}{r_{o}}\right)^{4}-1}-\frac{\epsilon}{r_{o}^{2}}\left[-1+f(r)+rf^{\prime}(r)\right]=0\ .$ (68)
So, it is convenient to introduce the nondimensional coordinate $\rho\equiv r/r_{o}>1$, and the nondimensional parameter
$\alpha\equiv\frac{\epsilon}{r_{o}^{2}}\ .$ (69)
Thus, the solution to Eq. (68) reads
$f(\rho)=1-\frac{C}{\rho}-\frac{\rho^{2}}{3\ \alpha}+\frac{\lambda}{3\alpha}\ \sqrt{\rho^{4}-1}+\frac{2\lambda}{3\alpha\ \rho}\,\ \Phi(\rho)\ .$ (70)
Here $C$ is an integration constant, and $\Phi(\rho)$ is
$\Phi(\rho)\equiv 2i\ \text{K}(-1)+\text{K}(-2)-i\ \text{F}(\arcsin\rho|-1)\ ,$ (71)
where $\text{F}(\varphi|m)$ is the elliptic integral of the first kind, and $\text{K}(m)$ is the complete elliptic integral of the first kind. $\Phi(\rho)$ is a real function in the domain $\rho>1$. The constants associated with K ensure that $\Phi(\infty)=0$; so $C$ is proportional to the mass $M$ of the black hole. Thus, the behavior of $f(r)$ at infinity is
$f(r)=1-\frac{2MG}{r}-\frac{\Lambda}{3}\ r^{2}+\frac{\beta^{2}(q^{2}+p^{2})}{2\lambda\epsilon}\ r^{-2}+\frac{\beta^{4}(q^{2}+p^{2})^{2}}{40\lambda^{3}\epsilon}\ r^{-6}+\mathcal{O}(r^{-10})\ .$ (72)
[…]
$\left(\frac{r}{r_{o}}\right)^{2}+\lambda\ \sqrt{\left(\frac{r}{r_{o}}\right)^{4}+1}+\frac{\epsilon}{r_{o}^{2}}\left[-1+f(r)+rf^{\prime}(r)\right]=0\ ,$ (73)
whose solution is
$f(\rho)=1-\frac{C}{\rho}+\frac{\rho^{2}}{3\ \alpha}-\frac{\lambda}{3\alpha}\ \sqrt{\rho^{4}+1}+\frac{2\lambda}{3\alpha\ \rho}\,\ \Psi(\rho)\ ,$ (74)
where
$\Psi(\rho)\equiv\sqrt{i}\left[2\ \text{K}(-1)-i\ \text{K}(2)+\ \text{F}(i\ \sinh^{-1}(\sqrt{i}\ \rho)|-1)\right]$ (75)
is a real function over the entire domain $\rho>0$ that satisfies $\Psi(\infty)=0$. The behavior of $f(r)$ at infinity is
$f(r)=1-\frac{2MG}{r}-\frac{\Lambda}{3}\ r^{2}+\frac{\beta^{2}(q^{2}+p^{2})}{2\lambda\epsilon}\ r^{-2}-\frac{\beta^{4}(q^{2}+p^{2})^{2}}{40\lambda^{3}\epsilon}\ r^{-6}+\mathcal{O}(r^{-10})\ .$ (76)
Comparing the metrics (72) and (76), we see that the choice of an imaginary parameter $\beta$ in the Lagrangian changes the metric just to the order $r^{-6}$. Compared with the Reissner-Nordström metric, which is a solution to the Einstein-Maxwell equations,
$f_{RN}(r)=1-\frac{2MG}{r}-\frac{\Lambda}{3}\ r^{2}+GQ^{2}\ r^{-2}\ ,$ (77)
we find that the role of $G$ is played by $\beta^{2}/(2\lambda\epsilon)$, as in Eq. (60). Thus, the parameter $\alpha$ defined in Eq. (69) can be rewritten as
$\alpha=\frac{\beta}{2G\sqrt{q^{2}+p^{2}}}\ .$ (78)
V.3 Metric horizons
In what follows, we will work with asymptotically flat geometries; so we will use $\lambda=1$ to make zero the cosmological constant $\Lambda$.
To locate the horizons of the metric, one must find the roots of the function $f(\rho)=g_{00}$. A graphic analysis of $f(\rho)$ in Eqs. (70) and (74) easily shows that there may be two, one or no horizons depending on the relation between the integration constant $C$ and the parameter $\alpha$. This is analogous to the case of Reissner-Nordström geometry, where the limit between one and two horizons is marked by the equality of the roots of $f$ when $GM^{2}=Q^{2}$ (extremal black hole). In a more general case, when roots cannot be easily computed, the extremal case can be obtained by looking for the conditions for the vanishing of $f$ and $f^{\prime}$ at the same point $\rho_{h}$. This double condition results in two equations to find $\rho_{h}$ and the relation $C=C(\alpha)$. For instance, in Reissner-Nordström one obtains
$1-\frac{2MG}{r_{h}}+\frac{GQ^{2}}{r_{h}^{2}}=0\ ,\ \ \ \ \frac{2MG}{r_{h}^{2}}-2\frac{GQ^{2}}{r_{h}^{3}}=0\ \ \ \ \Rightarrow\ \ \ r_{h}=GM\ ,\ \ \ \sqrt{G}M=Q\ ;$ (79)
at the level of nondimensional variables, this corresponds to $\rho_{h}=C/2$, $C=\sqrt{2/\alpha}$ (i.e., $\rho_{h}=1/\sqrt{2\alpha}$).
Applying this criterion to the function (70), it is found that
$\rho_{h}=\frac{\sqrt{1+\alpha^{2}}}{\sqrt{2\alpha}}\ .$ (80)
Note that $\rho_{h}\geq 1$, so the horizon of the extremal black hole always lies inside the domain of the coordinate $\rho$ ($\rho_{h}=1$ for $\alpha=1$). The function $\rho_{h}(\alpha)$ has two branches that span the range $1<\rho_{h}<\infty$. We are interested in the branch where $0<\alpha<1$, since the Reissner-Nordström horizon is recovered in the limit $\alpha\rightarrow 0$. The relation between $C$ and $\alpha$ allowing the black hole to be extremal with horizon at $\rho_{h}$ can be obtained from $f(\rho_{h})=0$; it yields
$C_{\text{extr}}=\frac{2\,\sqrt{1+\alpha^{2}}}{3\,\sqrt{2\alpha}}+\frac{2}{3\,\alpha\ }\,\ \Phi\left(\frac{\sqrt{1+\alpha^{2}}}{\sqrt{2\alpha}}\right)\ .$ (81)
Instead, if the determinantal Lagrangian is built with an imaginary value of $\beta$, then the geometry will be characterized by $f(\rho)$ in Eq. (74). By applying the same procedure, one obtains the horizon of the extremal black hole at
$\rho_{h}=\frac{\sqrt{1-\alpha^{2}}}{\sqrt{2\alpha}}\ ,$ (82)
while the relation between $C$ and $\alpha$ becomes
$C_{\text{extr}}=\frac{2\,\sqrt{1-\alpha^{2}}}{3\,\sqrt{2\alpha}}+\frac{2}{3\,\alpha\ }\,\ \Psi\left(\frac{\sqrt{1-\alpha^{2}}}{\sqrt{2\alpha}}\right)\ .$ (83)
In both cases the Reissner-Nordström limit is retrieved:
$C_{\text{extr}}=\sqrt{\frac{2}{\alpha}}+\mathcal{O}(\alpha^{3/2})\ .$ (84)
Figure 1 shows how the existence and number of horizons depend on the relation between the parameters $\alpha$ and $C$.
Figure 1: Plot of the function $f(\rho)$ when $C=C_{\text{extr}}$ (Eq. (70)) and when small deviations of that value are considered. It was chosen $\lambda=1$ and $\alpha=0.1$.
For $C$ (the parameter proportional to the black-hole mass) greater than $C_{\text{extr}}$ the curve exhibits two roots, which indicates the presence of two horizons. For $C<C_{\text{extr}}$, there are no horizons. A similar graphic is obtained for the solution (74) associated with an imaginary value of $\beta$ in the determinantal Lagrangian (although $\rho\in(0,\infty)$, and $f$ diverges at $\rho=0$).
V.4 The smoothing of singularities
As mentioned previously, the choice of an imaginary $\beta$ in the determinantal Lagrangian turns the picture $g$-electrodynamics into its standard Born-Infeld form. In fact, the electric field of the point charge becomes regular at the origin; thus, the field invariants $S$ and $P$ will not diverge at the origin as long as the magnetic monopole $p$ is zero. Then, it is natural to wonder about the impact of this regularity on the geometry of spacetime. First, note that the condition $p=0$ has no impact on the geometry because $p$ enters the geometry in the combination $(q^{2}+p^{2})^{1/2}$, since that is the way $q$ and $p$ appear in the energy-momentum tensor (65) as a factor in $r_{o}$. Second, the regular behavior of the electromagnetic field does not prevent the geometry from diverging. A direct inspection of Eq. (66) shows that the curvature scalar $R$ diverges regardless of whether $\beta$ is real or imaginary in the determinantal Lagrangian. If $\beta$ is real (anomalous case), then $R$ diverges at $\rho=1$ ($r=r_{o}$). If $\beta$ is imaginary (standard case), then $R$ is obtained from Eq. (66) by replacing $\beta^{2}$ with $\beta^{2}$ (that is, $r_{o}^{4}\rightarrow-r_{o}^{4}$) and $\epsilon$ with $\epsilon$; thus, $R$ becomes regular at $\rho=1$ but still diverges at $\rho=0$.
The behavior of the function $f(\rho)$ is not better in the standard case than in the anomalous case. In fact, $f(\rho)$ is regular at $\rho=1$ in the anomalous case (but its derivative diverges). Instead, $f(\rho)$ diverges at $\rho=0$ in the standard case. Thus, the way in which the Ricci tensor (67) diverges in each case is quite different. In the anomalous case, it results (we use $\lambda=1$ to exclude the contribution of the cosmological constant)
$R^{t}_{\ t}=R^{r}_{\ r}=\frac{1}{2\epsilon\,\sqrt{\rho-1}}-\frac{1}{\epsilon}+\mathcal{O}(\sqrt{\rho-1})\ ,\ \ \ \ \ \ R^{\theta}_{\ \theta}=R^{\phi}_{\ \phi}=-\frac{1}{\epsilon}+\mathcal{O}(\sqrt{\rho-1})\ ,$ (85)
while in the standard case it is
$R^{t}_{\ t}=R^{r}_{\ r}=\frac{1}{\epsilon}+\mathcal{O}(\rho^{2})\ ,\ \ \ \ \ \ R^{\theta}_{\ \theta}=R^{\phi}_{\ \phi}=-\frac{1}{\epsilon\,\rho^{2}}+\frac{1}{\epsilon}+\mathcal{O}(\rho^{2})\ .$ (86)
We can compare these results with the Ricci tensor in Reissner-Nordström geometry,
[…]
Therefore, Born-Infeld electrogravity, both in its anomalous and standard versions, alleviates the singular behavior of the geometry at $\rho=1$ and $\rho=0$ respectively. We can confirm this property at the level of the Riemann tensor by means of the Kretschmann scalar $K=R_{\alpha\beta\gamma\delta}R^{\alpha\beta\gamma\delta}$. Its expression in terms of $f(r)$ and its derivatives is given by
$K=\frac{[1-f(\rho)]^{2}+\rho^{2}\,f^{\prime}(\rho)^{2}+\rho^{4}\,f^{\prime\prime}(\rho)^{2}}{r_{o}^{4}\ \rho^{4}}\ .$ (88)
In the anomalous case (70), the behavior of $K$ at $\rho=1$ is
$K=\frac{1}{\epsilon^{2}\,(\rho-1)}-\frac{4}{3}\,\frac{1+3\,\alpha\,C-2\,\Phi(1)}{\epsilon^{2}\,\sqrt{\rho-1}}+\mathcal{O}(1)\ ;$ (89)
while in the standard case (74), the behavior of $K$ at $\rho=0$ is
$K=\frac{4}{3}\,\frac{(3\,\alpha\,C-2\,\Psi(0))^{2}}{\epsilon^{2}\,\rho^{6}}+\mathcal{O}(\rho^{-5})\ .$ (90)
On the other hand, the Kretschmann scalar for Reissner-Nordström geometry is
[…]
This means that the ability to smooth the geometry is better in the anomalous case (when $\beta$ is real in the determinantal Lagrangian), because the singularity has been shifted to $\rho=1$, on a sphere of area $4\pi\,r_{o}^{2}$; thus the pathological effect of the factor $\rho^{-4}$ in Eq. (88) is circumvented.
The divergent behavior of $K$ in the standard case of Eq. (90) is also better than in the Reissner-Nordström geometry, since $K$ keeps the behavior $\rho^{-6}$ characteristic of the (chargeless) Schwarzschild geometry. Noticeably, this behavior can still be improved by choosing a mass such that  [^9] This mass is extremal for $\alpha=1$ (see Eq. (83)).
$C=\frac{2\ \Psi(0)}{3\ \alpha}\ .$ (92)
This happens because this mass value regularizes the function $f(\rho)$ at the origin, since $f(\rho)$ in Eq. (74) turns to be
$f(\rho)=\frac{\alpha-1}{\alpha}+\frac{\rho^{2}}{3\,\alpha}+\mathcal{O}(\rho^{4})$ (93)
(a deficit angle is formed at the origin). With the choice (92), the Kretschmann scalar of the standard case (74) becomes
$K=\frac{1}{\epsilon^{2}\,\rho^{4}}-\frac{2}{3\epsilon^{2}\,\rho^{2}}+\frac{6}{5\epsilon^{2}}+\mathcal{O}(\rho^{2})\ .$ (94)
Instead, the behavior of $K$ in the anomalous case cannot be improved by choosing a particular value for $C$, since the first term in Eq. (89) does not depend on $C$. Neither the behavior of $K$ in Reissner-Nordström geometry could be improved in this way. Consistently, the same $r^{-4}$ dependence of the Kretschmann scalar was found in [28] solving the Einstein equations sourced by the Born-Infeld electromagnetism, for a dyonic black hole and a critical value of the mass.
Finally, the existence of a true singularity is confirmed by the fact that both $\rho=1$ in the anomalous case, or $\rho=0$ in the standard case can be reached in a finite proper time. This is because the behavior of $f$ at the singularity is either regular (anomalous case) or the same as in an ordinary black hole (standard case). Therefore, spacetime is geodesically incomplete.
VI Conclusions
We have analyzed a (determinantal) Born-Infeld-like theory in which the gravitational and electromagnetic sectors are coupled through a single determinant. The field equations were obtained by independent variations of the metric, the affine connection and the electromagnetic potential using the Palatini formalism. The metric variation gives a constraint equation (reflecting the absence of metric derivatives in the Lagrangian), while the dynamical content comes from the variations with respect to the connection and the potential. Working with the symmetric part of the Ricci tensor allowed us to set the torsion to zero, then the connection equation enforces metric compatibility and fixes the connection uniquely to the Levi–Civita one.
The field equations are naturally written in terms of the inverse of an auxiliary tensor $\bar{q}^{\mu\nu}$, so to express the dynamics in terms of the physical fields, we considered two complementary formulations. On the one hand, the theory reads more naturally as standard Born-Infeld electrodynamics on an effective background defined by the symmetric tensor $\mathcal{G}_{\mu\nu}=g_{\mu\nu}+\epsilon R_{(\mu\nu)}$. On the other hand, by combining the dynamical equations, one obtains an equivalent formulation in the physical metric $g_{\mu\nu}$; this second description displays a sign change in the electromagnetic invariants and therefore we referred to it as “anomalous”. We showed that the two pictures, $\mathcal{G}$” and “gg”, are related by a set of identities that map the pseudoscalars built with $\mathcal{G}$ into those built with $g$ and that also relate the corresponding volume elements (Eqs. (47) and (49)). In particular, these relations explain how the apparent coupling of the electromagnetic field to curvature in picture $\mathcal{G}$ (Eq. (37)) is reabsorbed and translated into a minimally coupled Born-Infeld structure in picture $g$, in agreement with the equivalence principle.
To explore the implications of the model on the singular behavior of the solutions, we analyzed spherically symmetric dyonic configurations. The electrodynamics in the anomalous case (real $\beta$) leads to an electric field that becomes singular at a finite radius $r_{o}$, restricting the domain of the solution to $r>r_{o}$, while the magnetic field keeps its Coulomb form. If instead the parameter $\beta$ is taken imaginary (which is allowed because only even powers of $\beta$ appear in the Lagrangian), the electric field takes the standard Born-Infeld form but the magnetic monopole is divergent at the origin. Remarkably, even though the electric and magnetic monopoles enter the field invariants $S$ and $P$ unequally, they contribute to the energy and pressure on an equal footing through the combination $\sqrt{q^{2}+p^{2}}$, a feature that is inherited by the geometry. The obtained geometric solution approaches the Reissner–Nordström form asymptotically (with an effective coupling replacing Newton’s constant), and the family of solutions admits two, one (extremal) or no horizons depending on the mass and parameter choices.
Concerning singularities, Born-Infeld electrogravity alleviates the geometric divergences compared to the Reissner–Nordström case, but nevertheless spacetime remains geodesically incomplete. The improvement is more pronounced in the (real $\beta$) anomalous case, in which the curvature singularity is shifted from the center to a sphere at $r=r_{o}$, mitigating the pathological inverse power behavior of the Kretschmann scalar. Notably, for imaginary $\beta$ there exists a solution for which the metric becomes regular at the origin and the singularity in the Kretschmann scalar is of order $r^{-4}$.
Finally, it would be interesting to investigate the cosmological solutions that arise from the Lagrangian considered in this work. In particular, in [9], the authors analyzed a type of BI gravity and found that the corresponding cosmological solution is free from singularities.
Acknowledgments: The authors were supported by Consejo Nacional de Investigaciones Científicas y Técnicas (CONICET) and Universidad de Buenos Aires (UBA).
Appendix A Determinant of $q$
The determinant of $q$ is the essential piece of Lagrangian (9). Since $q=\mathcal{G}+\beta~F=\mathcal{G}\ (I+\beta~\breve{F})$, then
$\det q=\det\mathcal{G~}\det(I+\beta~\breve{F})$ (95)
To compute $\det(I+\breve{F})$ we will use an expansion that is valid for any matrix $M$ in an arbitrary dimension $n$. Let us call $s_{k}$ the traces of the powers of $M$,
$s_{k}=\operatorname{Tr}(M^{k}),\quad 1\leq k\leq n\ .$ (96)
Then the determinant of IMI-M can be expressed as
$\det(I-M)=1+\sum_{k=1}^{n}p_{k},$ (97)
where the coefficients $p_{k}$ are related to the traces $s_{k}$ as follows
$\begin{split}p_{1}&=-s_{1},\\ p_{2}&=-\frac{1}{2}(s_{2}+p_{1}s_{1}),\\ p_{3}&=-\frac{1}{3}(s_{3}+p_{1}s_{2}+p_{2}s_{1}),\\ p_{4}&=-\frac{1}{4}(s_{4}+p_{1}s_{3}+p_{2}s_{2}+p_{3}s_{1}),\\$ $\vdots\end{split}$ (98)
In our case it is $n=4$, and $M=-\beta~\breve{F}$. Then,
$s_{1}=\operatorname{Tr}(M)=-\beta~\operatorname{Tr}(\breve{F})=-\beta~\operatorname{Tr}(\mathcal{\bar{G}}F)=0~,$ (99)
[…]
$s_{2}$ $=$ $\operatorname{Tr}(MM)=\beta^{2}~\breve{F}_{~~\lambda}^{\nu}\breve{F}_{~~\nu}^{\lambda}=-4\beta^{2}\breve{S}~,$ (100)
$s_{3}$ $=$ $\operatorname{Tr}(MMM)=-\beta^{3}~\breve{F}_{~~\lambda}^{\nu}\breve{F}_{~~\rho}^{\lambda}\breve{F}_{~~\nu}^{\rho}=0\ ,$ (101)
$s_{4}$ $=$ $\operatorname{Tr}(MMMM)=\beta^{4}~\breve{F}_{~~\lambda}^{\nu}\breve{F}_{~~\rho}^{\lambda}\breve{F}_{~~\eta}^{\rho}\breve{F}_{~~\nu}^{\eta}=\beta^{4}~\breve{F}_{~~\lambda}^{\nu}(-2\breve{S}~\delta_{\eta}^{\lambda}+^{\ast}\breve{F}_{~\rho}^{\lambda}~{}^{\ast}\breve{F}_{\eta}^{\rho})~\breve{F}_{~~\nu}^{\eta}$~ (102)
$=$ $\beta^{4}~(8~\breve{S}^{2}+\breve{P}^{2}~\delta_{\rho}^{\nu}\delta_{\nu}^{\rho})=4\beta^{4}~(2~\breve{S}^{2}+\breve{P}^{2})$
So it is
$p_{1}=0~,~~~~~p_{2}=-\frac{s_{2}}{2}=2\beta^{2}\breve{S}~,~~~~~p_{3}=0~,~~~~~p_{4}=-\frac{1}{4}(s_{4}+p_{2}s_{2})=-\beta^{4}~\breve{P}^{2}~.$ (103)
Therefore
$\det q=\det\mathcal{G~}\det(I+\beta~\breve{F})=(\det\mathcal{G)~}(1+2\beta^{2}\breve{S}-\beta^{4}~\breve{P}^{2})~.$ (104)
Let us mention that the same treatment could be given to $\det\mathcal{G}=\det g\,\det(\delta^{\mu}_{\nu}+\epsilon R^{\mu}_{\ \nu})$, or even to $\det q=\det g\,\det(\delta^{\mu}_{\nu}+\epsilon R^{\mu}_{\ \nu}+\beta F^{\mu}_{\ \nu})$. However, the traces of $R^{\mu}_{\ \nu}$ and its powers are all different from zero, which leads to a much more complicated expression.
Appendix B The projective mode
In Cartan language, curvature is expressed as a family of 2-forms,
$\widetilde{\widetilde{\mathcal{R}}}{}^{a}_{\;b}\equiv d\tilde{\omega}^{a}_{\;b}+\tilde{\omega}^{a}_{\;c}\wedge\tilde{\omega}^{c}_{\;b}\ ,$ (105)
labeled by two tangent space indices. The family of 1-forms $\tilde{\omega}^{a}_{\;b}$ constitutes the spin connection, defined as
$\tilde{\omega}^{a}_{\;b}=\Gamma^{a}_{\;bc}{\tilde{e}}^{c},$ (106)
where $\{\tilde{e}^{c}\}$ is the co-tangent space basis, dual to the tangent basis $\{\vec{e}_{c}\}$, and $\Gamma^{a}_{bc}$ are the components of the affine connection in that basis ($\nabla_{\vec{e}_{c}}\vec{e}_{b}=\Gamma^{a}_{bc}\,\vec{e}_{a}$). The usual components of the Riemann tensor $R^{a}{}_{bcd}$ then satisfy [^11] Coordinate bases are used in the main body of this article: $\{\tilde{e}^{c}\}\rightarrow\{dx^{\mu}\}$.
$\widetilde{\widetilde{\mathcal{R}}}{}^{a}_{\;b}=\frac{1}{2}\,R^{a}_{\,bcd}\;\tilde{e}^{c}\wedge\tilde{e}^{d}\ .$ (107)
The antisymmetry in the last two indices of $R^{a}{}_{bcd}$ reflects the 2-form nature of $\widetilde{\widetilde{\mathcal{R}}}{}^{a}_{\;b}$, while the indices $a$, $b$ label the different members of the family.
Let us now consider a projective transformation of the connection
$\tilde{\omega}^{a}_{\;b}\rightarrow\tilde{\omega}^{a}_{\;b}+\delta^{a}_{b}\;\tilde{\mathcal{A}}\$, (108)
where $\tilde{\mathcal{A}}=\mathcal{A}_{c}\,\tilde{e}^{c}$ is an arbitrary 1-form. This transformation is equivalent to [^12] Under this transformation, the covariant derivative of a vector $\bar{U}$ in the direction of a vector $\bar{V}$ acquires a term proportional to $\bar{U}$: $\nabla_{\bar{V}}\bar{U}\rightarrow\nabla_{\bar{V}}\bar{U}+\mathcal{A}_{c}V^{c}\,\bar{U}$. Along an autoparallel, where $\nabla_{\bar{U}}\bar{U}=0$, the term can be absorbed via a redefinition of the affine parameter.
$\Gamma^{a}_{bc}\longrightarrow\Gamma^{a}_{bc}+\delta^{a}_{b}\,\mathcal{A}_{c}\ .$ (109)
Under this transformation, the curvature becomes
[…]
which, using the antisymmetry of the wedge product, simplifies to
$\widetilde{\widetilde{\mathcal{R}}}{}^{a}_{\;b}\rightarrow\widetilde{\widetilde{\mathcal{R}}}{}^{a}_{\;b}+\delta^{a}_{b}\;d\tilde{\mathcal{A}}\$, (111)
or, in a coordinate basis,
$R^{\lambda}_{\;\rho\mu\nu}\rightarrow R^{\lambda}_{\;\rho\mu\nu}+\delta^{\lambda}_{\rho}\,(\partial_{\mu}{\mathcal{A}}_{\nu}-\partial_{\nu}{\mathcal{A}}_{\mu})\ .$ (112)
Consequently, the Ricci tensor transforms as
$R_{\rho\nu}\rightarrow R_{\rho\nu}+(\partial_{\rho}{\mathcal{A}}_{\nu}-\partial_{\nu}{\mathcal{A}}_{\rho})\ .$ (113)
Even if the Ricci tensor was initially symmetric, the projective transformation breaks that symmetry. However, in a Lagrangian of the form (9)–(10), a variation of the connection of the type $\delta\Gamma^{\alpha}_{\mu\nu}=\delta^{\alpha}_{\mu}\,\delta{\mathcal{A}}_{\nu}$ is indistinguishable from a variation of the electromagnetic potential $\delta A_{\nu}=\epsilon\beta^{-1}\delta\mathcal{A}_{\nu}$. This means that, in a Lagrangian like the one considered here, the projective degrees of freedom are degenerate with the electromagnetic ones. Freezing the projective variations of the connection (thus preserving the symmetry of the Ricci tensor) and retaining only variations of the electromagnetic potential is merely a gauge fixing that resolves this ambiguity.